The length of a rectangle is 5 times the width. The area is 720 square centimeters. Find the length and width of the rectangle.
Answer by addingup(3677)

The area of the rectangle is:
Area = Length*Width
The problem says that the length is 5 times the width:
L = 5W In the equation above let's substitute Length with this value:
A = 5W*W Or:
5w*W = A Now let's plug in the numbers:
5W*W = 720=
5*W*W= 720 Multiply W:
5*W^2= 720 Divide both sides by 5:
W^2= 144 Take the square root on both sides:
W= 12 This is the width, and the Length:
L= 5*12= 60
Proof:
12*60= 720 Multiply on left:
720= 720 We have the correct answer